OpenStudy (anonymous):

Oh my this is just too hard for me...how do i solve this.. x^5 - x + 1 = 0

OpenStudy (jiteshmeghwal9):

Is this \(x^2-x+1=0\) ?

OpenStudy (anonymous):

no \[x^{5} -x +1=0\]

OpenStudy (jiteshmeghwal9):

hmm.....

OpenStudy (anonymous):

a quintic equation

OpenStudy (jiteshmeghwal9):

\(x^5-x=-1\)

OpenStudy (anonymous):

nop quintic...

OpenStudy (jiteshmeghwal9):

sorry!

OpenStudy (anonymous):

and its so hard to solve i don know

OpenStudy (jiteshmeghwal9):

i m trying ok :)

OpenStudy (anonymous):

ok

OpenStudy (jiteshmeghwal9):

this quintic equation's root can't be expressed as radicals .

OpenStudy (hezekieli):

No idea for the answer but does it help to start with:\[x \left( x ^{4} - 1 \right) = 1\]?

OpenStudy (anonymous):

mmm...i think i can try from there...luckily i have the answer with me...

OpenStudy (hezekieli):

visually, the answer is somewhere ]-1.2;-1.1[

OpenStudy (anonymous):

I don't know if it will converge, you could try Newton's method: Make a guess, say -1.25 (x_0) then x_1 = x_0 - f(0)/f'(0) and repeat till u get close..

OpenStudy (anonymous):

yeah i tried Newton method and got x=-1.1679 or closer to this